In an experiment, a ball is released from a building. If the height of the ball after released is given by the equation h= -4.9t2+19.6t=686, where h is the height of the ball and t is the elapsed time after it is released, after how many seconds will the ball hit the ground?
Answer by ikleyn(52781)

.

            Your equation is written INCORRECTLY.

            The correct form is  h(t) = -4.9t^2 + 19.6t + 686 ,  where  h  is measured in feet.


The ball will hit the ground, when  h(t) = 0,  or, equivalently,


    -4.9t^2 + 19.6t + 686 = 0.


Simplify by dividing by (-4.9) to get


    t^2 - 4t - 140 = 0,


    (t-14)*(t+10) = 0.


Of the two roots,  t= 14  and  t= -10,  only positive t = 14 is meaningful, giving the


ANSWER.  t= 10 seconds.

Solved.